You are on page 1of 8

FEED BACK

PROJECT: ENNORE UNTT-4

No. 2
- 110 MW (RAM)

SUBJECT: HIGH AXIAL SHIFT IN 110 MW TURBINE

OBSERVATION

Unit No. 4 whi ch was runnin g at a lo ad of 40 MW wi th a high cu rtis stag e pressure of 62 KSC tripped at 16.01 Hrs. on 09.04.2001 on high axial shift (As indicated in the meter on that day, the rotor showed its movement towards generator). Simultaneously all other differential movements showed a rotor movement of 0.40 mm towards generator. These observations were made from the trend curve as taken at the time of tripping. Also from the trend curves (refer Curve No.1) taken on 08.04.2001 it was found that the axial shift which had gone to +0.5 mm (movements of rotor towards HP) was found to have slowly gone to -0.3 for which no reason could be immediately found out. The trends of TSI on 08.04.2001 showed that whenever a movement of rotor was there (may be due to load changes) as shown by the axial shift, the same was reflected on the DEs of HP, IP and LP. (Refer Trend Curves 2 & 3). ACTION SUGGESTED From the above observation, it was concluded that since there is movement of rotor by 0.4 mm as indicated by turbovisory instruments, the bearing No.2 may be opened and inspected. Simultaneously the axial shift pick up also may be calibrated. Before the trip, the HPDE instrument was showing high differential expansion readings. Hence it was suggested that the HPDE also may be checked for calibration availing the shutdown.

04
ACTION TAKEN Accordingly the bearing No. 2 was opened on 14.04.2001 after stopping barring gear. No abnormality was found on the bearing except for some rub marks on the thrust pads of the non-working side. These marks were removed by scraping and the bearing No.2 was assembled back. M/c was put back on baring gear and found OK. The axial shift pick up was taken out and was calibrated. The pick up was found to be OK except for the cable joint fault. This lug joint was replaced with soldered joint. Metallic hose was provided fully covering the pick up cable. The zero of the axial shift reading was adjusted at the standstill position of rotor after stopping barring gear motor on 17.04.2001. All the differential expansion pick ups were recalibrated and adjusted for zero with rotor in standstill condition after the metal temperatures of m/c come down below 500C.

ANALYSIS AND CONCLUSION OF THE OCCURRENCE - As per report by BHEL/SAS, the rotor had moved towards HP side when the machine was on load and was normally working on the non-working pad side. This movement was due to high curtis stage pressure. The same was indicated by TSI also during normal working of the Unit which was indicated as 0.5 mm as against a designed value of 0.1/0.2 mm. - While calibrating and inspecting the axial shift pick up, the following observations are made: a) The cable joint which was provided by providing straight lugs had got oxidized. b) The PVC insulation tape and the Teflon tape wound on the joint had got burnt. The reasons for the above cable failure is attributable to the high temperature steam impingement on 20.02.2001 when the flange in the W pipe of MP inlet failed. The steam impingement caused a reduction in life of the joint. This was acting on the joint as slow poison and weakened the joint gradually.

05
This weak cable joint caused voltage drop and an error in the axial shift reading on 07.04.2001 and the reading slowly drifted from +0.5 to -0.3. The machine was working with this drift till 09.04.2001 when another drift of 0.3 mm occurred suddenly and caused the tripping of the Unit. (Refer Curve No.2) at the set value of - 0.6 mm. In fact, the rotor was actually at 0.4 mm position due to high curtis stage pressure. But returned to its neutral position on tripping. Since only an axial shift indication is used for protection also system reliability problem experienced. TEETHING PROBLEM : The turbovisory panel, supplied by M/s Global Weighing Ltd (earlier known as Philips) through EDN, had reached site without the completion of internal wiring of the cards and instruments. These works were carried out at site with short duration of time since the supplier engineer who reached site also was not aware of the panel condition. The unit had teething problem in the TSI system and Unit had tripped on axial shift/HP DE due to malfunctioning cable joints and cards. The defective cable connectors were replaced with straight lug joints at site in consultation with supplier. Hence site suggested two out of three logics protection system during November, 2000. RECOMMENDATION 1. As per the modem control concept, all the trips are to be provided through 2/3 logic. The same is suggested for axial shift also. This logic is provided for axial shift at Jindal (2x130 MW sets) by BHEL/Hyd. If at all any cable joint is to be done in this measuring circuit, it should be a soldered joint only to avoid leakage of joint.

2.

06

07

08

09

10